Solve by the elimination method.
9x-9y=36
7y-3x= -14
What is the solution of the system?
Can someone explain this to me? I do not understand this at all...not of it.

Answer by scott8148(6628)

in order to ELIMINATE one of the variables, you need to get the absolute value of the coefficients to match
__ then you can either add or subtract the equations and the variable will be cancelled (eliminated)

multiplying the second equation by 3 __ 21y-9x=-42
__ adding to first equation __ 9x-9y+21y-9x=36-42 __ 12y=-6 __ y=-1/2

substituting __ 7(-1/2)-3x=-14 __ -3x=-21/2 __ x=7/2